AS annouched in local meetings and during the coll meeting the clean-up of hardware system of trigger has started. It is still work in progress. All calbles & modules for which the function is not either needed or unknown will be taken out and the layout better organized. The sync and pulser trigger has been re-established in working order i.e. T7 and T8. Please do not add stuf without consulting me. Special note: The mult-tiles stuff has also been moved and can/will be setup again when YK comes Monday for the gain matching of tiles. Flemming ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov
This archive was generated by hypermail 2b29 : Fri May 04 2001 - 17:03:48 EDT